# Coupling without Communication and Drafter-Invariant Speculative Decoding

### v2

@@ 5 Application: Drafter-Invariant Speculative Decoding
The end result is that Alice’s predictions are accurate with probability exactly equal to $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. However, this comes with a potential caveat: the output of the process, depends on both $\mathcal{P}$ and $\mathcal{Q}$. Hence, if the small “drafter” model changes, the large model output also changes. This is undesirable: as discussed in Section 1, the output of an LM is ideally fixed given a fixed random seed, no matter what optimizations are used to accelerate inference of the network. This property allows for users to easily reproduce results and facilitates easier unit testing and debugging. Naively, speculative sampling destroys this property.
However, the issue is easily fixed: if, instead of an optimal coupling, we use a communication-free protocol to sample from the large model, then necessarily the models output is independent of the drafter model. We call the resulting approach “Drafter-Invariant Speculative Decoding”. There is a price to pay for drafter-invariance: by using a communication-free protocol, we can only ensure that the draft tokens are correct with probability $c=\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ instead of $c=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Since the expected number of sequential tokens drafted correctly is $\frac{1}{1-c}$, this could lead to a roughly 2×2$\times$2 × reduction. Fortunately, as we will see shortly, the no-communication protocols analyzed in Section 3, especially our Gumbel sampling approach, tend to outperform the worst-case bound of Theorem 3 for real-world distributions arising from language models.
@@ 5.1 Experimental Evaluation
As we can see in the figure, Gumbel sampling achieves a higher collision probability than Weighted MinHash, across all queries and tokens. Both methods typically outperform the worst-case bound of $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ proven in Theorem 3, often nearly matching the best possible collision probability of $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$.
@@ 6 Coupling with Low-Communication
To prove Theorem 4, we first introduce an “idealized” communication protocol, Protocol 4, that proceeds in rounds. At each round, Alice and Bob communicate at most one index consisting of $O(\log n)$ bits, and one probability, i.e., a real-valued number in the interval $[0,1]$. We show that this protocol terminates with an expected constant number of steps. We then show how to appropriately discretize Alice and Bob’s distributions in a pre-processing step so that the probability can be communicated with $O(\log(n/\epsilon))$ bits. The same discretization strategy can be used to give a concrete communication complexity bound for Protocol 1, which naively requires communicating real-valued probabilities. The result would be a protocol requiring $O(n\log(n/\epsilon))$ bits of communication. Theorem 4 offers and exponential improvement on this baseline.
@@
6:         Send approve to Bob if $p_{j}<u_{k}-w$, send reject otherwise.
@@
Protocol 4 implements this “post-rejection” step in a more communication efficient way by using a dart-throwing approach similar to Weighted MinHash. Specifically, Bob attempts to sample from regions of the number line where $q_{j}>p_{j}$, as illustrated by the hatched regions in Figure 3. A uniform sample from these regions amounts to a sample from $\mathcal{Q}^{\prime}$. To obtain such a sample, Bob throws a dart and, if it lands in the region $[j-1,j-1+q_{j}]$ for some $j$, he sends the dart to Alice to verify that it also lies above $j-1+p_{j}$. If it does, then the dart has landed in the hatched region, so they terminate the process. If not, Bob tries again. We prove that the number of rounds required for the process to terminate is constant in expectation:
@@
Furthermore, the expected number of messages before the protocol terminates is $2$.
@@
The bottom of the fraction above is exactly the probability that Bob rejects at Line 6, so overall, $\Pr[b=j]=\Pr[b=j\mid b=a]+\Pr[b=j\mid b\neq a]=\min(q_{j},p_{j})+q_{j}-\min(q_{j},p_{j})=q_{j}$, as desired.
@@
$\mathbb{E}[c]=\Pr[a=b]\cdot 1+\Pr[a\neq b]\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}=1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}=2.$
@@ Acknowledgement
This work was partially support by NSF Grant #$\#2106888$.

### v4

@@ 5 Application: Drafter-Invariant Speculative Decoding
The end result is that Alice’s predictions are accurate with probability exactly equal to $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. However, this comes with a potential caveat: the output of the process depends on both $\mathcal{P}$ and $\mathcal{Q}$. Hence, if the small “drafter” model changes, while the token distribution never changes, the exact token sampled by the large model very well could. This is undesirable: as discussed in Section 1, the output of an LM is ideally fixed given a fixed random seed, no matter what optimizations are used to accelerate inference of the network. This property allows for users to easily reproduce results and facilitates easier unit testing and debugging. Naively, speculative sampling destroys this property.
Fortunately, this issue can be fixed with a communication-free protocol! If we use a communication-free protocol to sample from the large model, then necessarily the models output is independent of the drafter model. We call the resulting approach “Drafter-Invariant Speculative Decoding”. There is a price to pay for drafter-invariance: by using a communication-free protocol, we can only ensure that the draft tokens are correct with probability $c=\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ instead of $c=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Since the expected number of sequential tokens drafted correctly is $\frac{1}{1-c}$, this could lead to a roughly $2\times$ reduction, although as we will see, the communication-free protocols analyzed in this work, especially our Gumbel sampling approach, tend to outperform the worst-case bound of Theorem 2 for real-world distributions arising from language models.
@@ 5.1 Experimental Evaluation
As we can see in the figure, Gumbel sampling achieves a higher collision probability than Weighted MinHash, across all queries and tokens, which is consistent with our Theorem 3. Both methods typically outperform the worst-case bound of $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ proven in Theorem 2, often nearly matching the best possible collision probability of $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$.
@@ 6 Coupling with Low-Communication
To prove Theorem 4, we first introduce an “idealized” communication protocol, Protocol 4, that proceeds in rounds. At each round, Alice and Bob communicate at most one index consisting of $O(\log n)$ bits, and one probability, i.e., a real-valued number in the interval $[0,1]$. We show that this protocol terminates with an expected constant number of rounds. We then show how to appropriately discretize Alice and Bob’s distributions in a pre-processing step so that the probability can be communicated with $O(\log(n/\epsilon))$ bits. The same discretization strategy can be used to give a concrete communication complexity bound for Protocol 1, which naively requires communicating real-valued probabilities. The result would be a protocol requiring $O(n\log(n/\epsilon))$ bits of communication. Theorem 4 offers and exponential improvement on this baseline.
@@
6:   If $p_{j}\geq u_{k}-w$, send reject to Bob. Otherwise send approve and break the loop.
@@
Protocol 4 implements this “post-rejection” step in a more communication efficient way by using a dart-throwing approach similar to Weighted MinHash. Specifically, Bob attempts to sample from regions of the number line where $q_{j}>p_{j}$, as illustrated by the hatched regions in Figure 3. A uniform sample from these regions amounts to a sample from $\mathcal{Q}^{\prime}$. To obtain such a sample, Bob throws a dart and, if it lands in the region $j-1,j-1+q_{j}]$ for some $j$, he sends the dart to Alice to verify that it also lies above $j-1+p_{j}$. If it does, then the dart has landed in the hatched region, so they terminate the process. If not, Bob tries again. We prove that the number of rounds required for the process to terminate is constant in expectation:
@@
Furthermore, the expected number of messages before the protocol terminates is at most $2$.
@@
The bottom of the fraction above is exactly the probability that Bob rejects at Line 6, so overall, $\Pr[b=j]=\Pr[b=j\land b=a]+\Pr[b=j\land b\neq a]=\min(q_{j},p_{j})+q_{j}-\min(q_{j},p_{j})=q_{j}$, as desired.
@@
$\mathbb{E}[c]=\Pr[a=b]\cdot 1+\Pr[a\neq b]\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\leq 1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\leq 2.$
@@ Acknowledgement
This work was partially support by NSF Grant #2106888\#2106888. Authors thank Mert Cemri, Nived Rajaraman, and Cheuk Ting Li for pointing out related works.